An environmental scientist wants to estimate the number of carp in a pond. He captures 100 carp, tags all of them, and releases them. A week later he captures 85 carp and records how many have tags. His results are shown in the table. Write and solve a proportion to estimate the number of carp in the pond.
Table...
Status.........Number Captured
Tagged 20
Not tagged 65

Answer by checkley77(12844)

100/x=20/65
20x=6,500
x=6,500/20
x=325 is the estimated number of carp in the pond.